# SoK: On Finding Common Ground in Loss Landscapes Using Deep Model Merging Techniques

### v1

@@ III-B1 Mode Connectivity
$\mathcal{L}_{p}=\max_{\lambda\in[0,1]}\mathcal{L}(p(\theta_{a},\theta_{b},\lambda))-\frac{1}{2}(\mathcal{L}(\theta_{a})+\mathcal{L}(\theta_{b}))$ (1)
@@
Two models are said to exhibit Mode Connectivity if there exists a path $p$ between them where the loss barrier is approximately zero [46, 47, 19]. Linear Mode Connectivity (LMC) [48, 47] is a stronger condition asserting that two modes can be connected by a straight line—of special interest as linearly connected solutions may lie in the same convex basin. Mode-connecting paths reveal the structure of the loss landscape and allow us to find diverse but equally performant solutions between two known models [49].
@@
In practice, we can discover paths of low loss between distinct neural networks via optimization (e.g., by parameterizing a Bézier curve) [50, 51, 46, 52, 53]. Surprisingly, solutions are often connected by simple curves, and even linear mode connectivity seems commonplace [50, 47]. Recent work has exploited this fact to ensemble networks that are close in the objective landscape, indicating that despite lying close together, sufficient diversity exists between these solutions to provide an advantage from ensembling [50]. Some empirical evidence suggests that geodesics between neural networks in the distribution space induced by the Fisher-Rao metric may correspond to paths between their parameters in objective space [52]. This finding links the notions of mode connectivity and interpolating between distributions. Additionally, it has been shown that linear interpolation between the parameters of two linearly connected models corresponds to linear interpolation between their activation maps [39]. This property holds even when the endpoints of interpolation are two networks fine-tuned on distinct tasks, provided that they stem from the same pre-trained checkpoint (see Section III-B3) [54].
III-B2 Permutation Symmetry
@@
Neural networks inherently have components that are invariant to permutations. For example, in a single linear layer of neurons, any permutation in the order of neurons would still result in the same activations for that layer. Therefore, we can consider any neural network to be a member of an equivalence class of other neural networks whose units are permuted but whose activations are the same. This permutation-invariance property has led some to conjecture that all solutions to Stochastic Gradient Descent (SGD) can be linearly connected if the natural permutation invariance of neural networks is accounted for [16]. This conjecture has spurred further investigation into the permutation symmetries inherent in the layers of neural networks and attempts to resolve them to achieve LMC between different solutions.
@@
Several theoretical results demonstrate that the objective landscapes of over-parameterized models are dominated by permutation-equivalent objective modes and that valleys in the loss landscape can be connected through permutation points [55, 56]. Additionally, many merging works provide substantial empirical evidence that aligning neurons by learning appropriate permutations in their order reveals linearly connected paths between models (see Section VI) [19, 20, 57, 51, 58]. This property holds even when models are trained on slightly differing data distributions [19, 20]. It has been observed that models trained on similar underlying distributions tend to have a similar distribution of singular values but may differ in their corresponding singular directions. This observation has led some to propose that the success of neuron alignment methods can be explained by the resulting alignment of dominant singular directions in the weight matrices by learned permutations of the order of neurons [59]. This observation suggests we can meaningfully associate task-specific performance with just a few singular directions in parameter space [59, 60].
III-B3 Training Trajectories
@@
The position of a neural network in the objective landscape is determined by the interaction between loss landscape geometry and gradient descent. Therefore, understanding training trajectories is paramount in model merging. A training trajectory is a sequence of points, each representing the parameters of a neural network, that evolves under some optimization algorithm, such as SGD, during the training process. Numerous efforts have been made to characterize how training proceeds from distinct initialization points under SGD and how this relates to the underlying objective landscape. Empirically, researchers have observed that SGD is initially unstable but soon reaches a “stable point”. After this point, subsequent fine-tuning causes networks to converge to the same loss basin even under different forms of SGD noise (e.g., batch order) [61, 47]. Some work has also demonstrated the converse, that deep neural networks with different random initializations learn functions with distinct properties and lie in distinct regions [62]. These works all support the conclusion that neural networks that share a significant portion of their training trajectory tend to lie in the same linearly connected region of the objective landscape. Conversely, those that are initialized separately lie in distinct regions of the objective landscape. A corollary of this finding is that models in the same objective basin generally do not require any form of neuron alignment for merging [21].
IV Ensembling
Ensembling is the most widely applied and reliable method to combine knowledge between models [63]. Ensemble methods are precursors to modern model merging techniques. In this section, we review the origins of ensembling and several prominent ensembling methods to frame our subsequent discussion of model merging techniques.
IV-A Overview
@@
Ensembling combines the predictions (as opposed to parameters) from several trained models to enhance accuracy at inference-time. Ensembles compute a weighted combination of output predictions from each model in the ensemble to form a prediction. The success of ensemble methods is often attributed to a reduction in model bias and variance that occurs when predictions are averaged from several statistical models [62, 63]. Ensembling is closely related to Bayesian Model Averaging techniques [64]. In classic Bayesian Model Averaging, the final prediction is a weighted average of the predictions of each of our models, weighted by the posterior probability of each model [65].
@@
where $y$ is the data label, $x$ the input data, $D$ the training data, and $\theta$ the neural network parameters. Integrating over all weight settings for a neural network is intractable; however, many works observe that neural networks consistently converge to solutions lying in large, flat basins in the loss landscape [64, 66, 67, 68, 69]. One explanation for this behavior is that these flat areas occupy large volumes in parameter space, and thus, stochastic gradient descent is exponentially more likely to land in these regions [66]. From a Bayesian perspective, these large, flat areas in parameter space have high probability density, $p\left(\theta\mid D\right)$, in the integral above. Therefore, an ensemble where we take the average predictions of several trained neural networks [70] can be thought of as an approximation of a Bayesian calculation of the posterior probability of $y$, with the integral estimated by the normalized summation of a few high probability parameter settings.
IV-B Ensemble Techniques
@@
Despite the reliability of ensemble methods, they can be impractical because they require training and deploying several deep neural networks simultaneously. Some methods attempt to reduce the cost of obtaining several models to be used as ensemble members. Snapshot Ensembles cycle through learning rates during training to produce several diverse models at various minima within a single training run [71]. Another approach is to exploit the local loss geometry around an existing model to discover useful ensemble members; for example, sampling new models from within the same objective basin as an existing model—as done in Fast Geometric Ensembling (FGE) [50]. This idea exploits the linear connectivity within an objective basin but has since been extended to exploit more complex forms of Mode Connectivity in the loss landscape. Benton et al. [72] develop a method that finds whole simplices of low loss in the objective landscape that connect several potential solutions. This method closely approximates a classic Bayesian model average where the posterior is uniformly distributed over a simplex.
@@
In addition to algorithmic advancements, recent increases in the number of freely available models have also reduced the cost of constructing ensembles. More pressing is the issue of runtime and memory efficiency, each of which scales linearly with the number of models in the ensemble. Mixture-of-Experts (MoE) routines reduce the inference cost associated with ensembling to that of a single model [73]. Rather than averaging the predictions of each ensemble member, these methods learn to compose components—at the granularity of whole networks, layers, or individual neurons—from each model to compute the final output prediction. This amounts to learning how to construct a new neural network for each input presented. Evolutionary Model Merging is a similar technique that uses an evolutionary algorithm to compose layers from each ensemble member to construct a new network [74].
IV-C Limitations of Ensembling
@@
Ensembles of several models have become much more straightforward to obtain thanks to advancements in our understanding of the loss landscape [50, 72, 71]. Still, ensemble techniques suffer from poor scaling in terms of runtime and memory costs, as we must use multiple models for inference. MoE methods are a boon for use cases where inference cost is a limiting factor. Unfortunately, MoE methods are not necessarily a universal solution. While the effective parameter count for a given input in a MoE model is lower, the total number of parameters in the mixture may be the same or even greater [75]. As individual “experts” in the mixture specialize in just a few prediction domains, we must train more parameters in total to achieve strong performance across all domains. Consequently, training and serving MoE models requires constantly exchanging large numbers of parameters from memory. This requires a significant investment in complex infrastructure to serve models of billions or trillions of parameters. MoE models also exhibit decreased sample efficiency during training; each parameter is not trained during every batch, increasing training costs. Though these methods offer exciting avenues for scaling and deploying ensemble models more efficiently, model merging offers the potential for equivalent performance at a significant cost reduction.
@@
Figure 2: Mode Convexity: Within an objective basin, one can linearly interpolate between known solutions to discover equally performant models with diverse behaviors.
V Weight Aggregation
@@
The fundamental technique underlying most model merging methods is weight aggregation. Given two or more models of the same architecture, one can compute an average or sum over the values for each corresponding parameter to obtain a set of new parameter values that comprise the merged model. These techniques are predicated on the principles of Bayesian Model Averaging, seeking to reduce estimation error from individual model bias and variance by combining several models for inference [76]. Though weight aggregation methods show promise compared to ensemble methods, they also introduce several challenges. Weight aggregation methods usually introduce additional hyperparameters that must be set via trial-and-error or by some heuristic method. Additionally, their failure modes are not fully understood. While some characterizations of their failure cases can be extrapolated from the literature (see Section VI below), more context on loss landscape geometry is necessary to fully understand their modes of failure. Below, we discuss several modes of weight aggregation, including simple averaging, weight steering via task arithmetic, and heuristic weighted averaging.
V-A Simple Averaging
@@
The most straightforward method of merging model parameters is to take a simple average over each corresponding parameter between two or more models, see Eq. (missing) 3. This technique is also employed in federated settings as FedSGD [27]. However, simple averaging does not consider each model’s relevance to the task at hand. This method also consistently fails if the training distributions or the training trajectories between the two models have diverged [20, 21, 19]. Despite this limitation, simple averaging is particularly useful in the “pre-train and fine-tune” model deployment paradigm, where models naturally share an architecture and are predisposed towards similar representations through pre-training [39, 54, 77]. LLMs are a good example of this kind of deployment.
$\theta^{\star}=\frac{1}{N}\sum_{i}\theta_{i}$ (3)
@@
A host of methods attempt to improve performance by using variations of simple averaging. Stochastic Weight Averaging (SWA) [64] averages checkpoints from a single model at different points in its training trajectory, demonstrating comparable performance to FGE [50]. Model Soups [78] are simple averages of fine-tuned models of the same architecture trained with different hyperparameters. Model soups often enhance Out-Of-Distribution (OOD) and zero-shot performance on downstream tasks. Sparse soups [79] reduce the inference costs associated with model soup methods by enforcing a sparsity constraint via an iterative prune-retrain fine-tuning technique similar to Iterative Magnitude Pruning (IMP) [80]. The similar SWAMP [81] method iteratively performs IMP and simple weight averaging between multiple fine-tuned models to produce a final sparse model. It has been observed that models closer to the center of a group of fine-tuned models stemming from the same pre-trained model display better OOD performance and robustness [82, 83].
Recent work has also demonstrated that several networks can be trained in tandem such that they form a linearly connected subspace of solutions. The midpoint of this subspace approaches the accuracy of an ensemble of independently trained models and demonstrates robustness to label noise [84]. Model Stock takes advantage of this observation by estimating a “central” network using only two fine-tuned models, significantly reducing the training effort required to construct model soups [83].
V-B Weight Steering
@@
Task Arithmetic [85] is a technique for manipulating model weights to enhance task-specific and OOD performance. A task vector—defined as the difference between the pre-trained and fine-tuned model weights for a task, see Eq. (missing) 4—can be thought of as the direction in weight space that the pre-trained model must travel to achieve better accuracy on that task.
$\Delta_{\text{task}}=\theta_{\text{task}}-\theta_{\text{pre}}$ (4)
@@
Interestingly, combining task vectors via simple arithmetic operations (addition and subtraction) provides predictable control over model behavior. For example, subtracting a task vector from the pre-trained weights reduces performance, while adding the same task vector improves performance on the corresponding task; adding two task vectors for distinct tasks to the pre-trained model weights enhances the resulting model’s performance on both tasks. This property is exploited by Model Breadcrumbs [86], a method that takes task vectors from several fine-tuned models, masks outlier directions, and combines them via addition into a single parameter update for the pre-trained checkpoint. Steering weights using task vectors has been observed to enable good performance in multiple tasks without degrading performance in individual tasks significantly, particularly for larger models [87]. TIES-merging [88] and DARE [89] improve upon Task Arithmetic by eliminating redundancies between task vectors, resulting in improved performance—particularly in the multi-task regime.
It is worth noting the connection between these methods and parameter-efficient fine-tuning schemes such as LoRA [90, 91, 92] and IA3 [93]. LoRA trains low-rank decompositions of weight matrices during fine-tuning that are added to the model weights for inference, while IA3 trains per-task scaling vectors in each layer that rescale the activations in that layer—equivalent to rescaling the weight matrix along corresponding directions in weight space. It is common to “swap” fine-tuning weights for different tasks by adding and subtracting out LoRA weights—a procedure analogous to how we use task vectors. LoReFT [94], inspired by LoRA, uses Distributed Alignment Search (DAS) [95, 96] to find task-specific directions in which to update low-rank fine-tuning weights.
@@
The success of weight steering methods implies that, within a single objective basin, we can meaningfully assign semantic properties to directions in weight space. Furthermore, we can leverage the linearly connected nature of local minima to smoothly interpolate between task-optimal parameter settings within a basin—even for multiple tasks. This raises the question: can we improve the performance of model merging techniques by calibrating the contribution of task-specific parameters to the final merged model?
@@
Figure 3: RegMean finds a new set of merged parameters that closely approximates the activation maps of the given source models in a manner analogous to least-squares regression.
V-C Heuristic Weightings

### v2

@@ 2.2.1. Mode Connectivity
(1) $\mathcal{L}_{p}=\max_{\lambda\in[0,1]}\mathcal{L}(p(\theta_{a},\theta_{b},\lambda))-\frac{1}{2}(\mathcal{L}(\theta_{a})+\mathcal{L}(\theta_{b}))$
@@
Two models are said to exhibit Mode Connectivity if there exists a path $p$ between them where the loss barrier is approximately zero (28; 40; 3). Linear Mode Connectivity (LMC) (95; 40) is a stronger condition asserting that two modes can be connected by a straight line. Linear connectivity is of special interest as linearly connected solutions may lie in the same convex objective basin. Mode-connecting paths reveal the geometric structure of the loss landscape and allow us to find diverse but equally performant solutions between two known models (66).
@@
In practice, we can discover paths of low loss between distinct neural networks via optimization (e.g., by parameterizing a Bézier curve) (41; 115; 28; 111; 100). Surprisingly, solutions are often connected by simple curves, and even linear mode connectivity is commonplace (41; 40). Recent work has exploited this fact to ensemble networks that are close in the objective landscape, indicating that despite lying close together, sufficient diversity exists between these solutions to provide an advantage from ensembling (41). Some empirical evidence suggests that geodesics between neural networks in the distribution space induced by the Fisher-Rao metric may correspond to paths between their parameters in objective space (111). This finding links the notions of mode connectivity and interpolating between distributions. Additionally, it has been shown that linear interpolation between the parameters of two linearly connected models corresponds to linear interpolation between their activation maps (142). This property holds even when the endpoints of interpolation are two networks fine-tuned on distinct tasks, provided that they stem from the same pre-trained checkpoint (see Section 2.2.3) (141).
Figure 2. Permutation Symmetry: An illustration of symmetry in dense linear layers. Changing the order of two neurons does not change the resulting activations.
2.2.2. Permutation Symmetry
@@
Neural networks inherently have components that are invariant to permutations. For example, in a single (dense) linear layer of neurons, any permutation in the order of neurons would still result in the same activations for that layer (see Fig. 2). Therefore, we can consider any neural network to be a member of an equivalence class of other neural networks whose units are permuted but whose activations are the same. This permutation-invariance property has led some to conjecture that all solutions to SGD can be linearly connected if the natural permutation invariance of neural networks is accounted for (30). This conjecture has spurred further investigation into the permutation symmetries inherent in the layers of neural networks and attempts to resolve them to achieve LMC between different solutions.
@@
Several theoretical results demonstrate that the objective landscapes of over-parameterized models are dominated by permutation-equivalent objective modes and that valleys in the loss landscape can be connected through permutation points (106; 12). Additionally, many merging works provide substantial empirical evidence that aligning neurons by learning appropriate permutations in their order reveals linearly connected paths between models (see Section 5) (3; 107; 81; 115; 120). This property holds even when models are trained on slightly differing data distributions (3; 107). Models trained on similar underlying distributions have been observed to learn weight matrices with similar distributions of singular values but differing singular directions. This observation has led some to propose that the success of neuron alignment methods can be explained by the resulting alignment of dominant singular directions in the weight matrices by learned permutations of the order of neurons (see Section 5.3) (61). This observation suggests we can meaningfully associate task-specific performance with just a few singular directions in parameter space (61; 117).
2.2.3. Training Trajectories
@@
The position of a neural network in the objective landscape is determined by the interaction between loss landscape geometry and gradient descent. Therefore, understanding training trajectories is paramount in model merging. A training trajectory is a sequence of points, each representing the parameters of a neural network, that evolves under some optimization algorithm, such as SGD, during the training process. Numerous efforts have been made to characterize how training proceeds from distinct initialization points under SGD and how this relates to the underlying objective landscape. Empirically, researchers have observed that SGD is initially unstable but soon reaches a “stable point”. After this point, subsequent fine-tuning causes networks to converge to the same loss basin even under different forms of SGD noise (e.g., batch order) (96; 40). Some work has also demonstrated the converse, that deep neural networks with different random initializations learn functions with distinct properties and lie in distinct regions (36). These works all support the conclusion that neural networks that share a significant portion of their training trajectory tend to lie in the same linearly connected region of the objective landscape (see Section 6.1.2). Conversely, those that are initialized separately tend to lie in distinct regions of the objective landscape. A corollary of this finding is that models in the same objective basin generally do not require any form of neuron alignment for merging (64).
3. Ensembling
Ensembling is the most widely applied and reliable method to combine knowledge between models (94). Ensemble techniques combine the output logits of several models or compose their layers rather than combining their parameters (using, say, an average) to produce a single merged model. This means that rather than producing one model from several, these techniques apply several models to a single inference task; this makes their deployment in the contemporary billion or trillion parameter regime difficult. Despite the fact that ensemble techniques are not true model merging techniques in the way that we define them (in the sense that they do not actually combine model parameters), ensemble techniques are precursors to much of the model merging literature, therefore, we review several prominent ensembling techniques below to frame our subsequent discussion of weight aggregation and neuron alignment.
3.1. Overview
@@
Ensembling combines the predictions (as opposed to parameters) from several trained models to enhance accuracy at inference-time. Ensembles compute a weighted combination of output predictions from each model in the ensemble to form a prediction. The success of ensemble methods is often attributed to a reduction in model bias and variance that occurs when predictions are averaged from several statistical models (36; 94). Ensembling is closely related to Bayesian Model Averaging techniques (62). In classic Bayesian Model Averaging, the final prediction is a weighted average of the predictions of each of our models, weighted by the posterior probability of each model (50).
@@
where $y$ is the data label, $x$ the input data, $D$ the training data, and $\theta$ the neural network parameters. Integrating over all weight settings for a neural network is intractable; however, many works observe that neural networks consistently converge to solutions lying in large, flat basins in the loss landscape (62; 54; 17; 49; 68). One explanation for this behavior is that these flat areas occupy large volumes in parameter space, and thus, stochastic gradient descent is exponentially more likely to land in these regions (54). From a Bayesian perspective, these large, flat areas in parameter space have high probability density, $p\left(\theta\mid D\right)$, in the integral above. Therefore, an ensemble where we take the average predictions of several trained neural networks (73) can be thought of as an approximation of a Bayesian calculation of the posterior probability of $y$, where one estimates the integral above by summation over a few parameter settings with high probability under the posterior.
3.2. Ensemble Techniques
@@
Despite the reliability of ensemble methods, they can be impractical because they require training and deploying several deep neural networks simultaneously. Some methods attempt to reduce the cost of obtaining several models to be used as ensemble members. Snapshot Ensembles cycle through learning rates during training to produce several diverse models at various minima within a single training run (52). Another approach is to exploit the local loss geometry around an existing model to discover useful ensemble members; for example, Fast Geometric Ensembling (FGE) (41) samples new models from within the same objective basin as an existing model. This idea exploits the linear connectivity within an objective basin but has since been extended to exploit more complex forms of Mode Connectivity in the loss landscape. Benton et al. (9) develop a method that finds whole simplices of low loss in the objective landscape that connect several potential solutions. This method closely approximates a classic Bayesian model average where the posterior is uniformly distributed over a simplex.
@@
In addition to algorithmic advancements, recent increases in the number of freely available models have also reduced the cost of constructing ensembles. More pressing is the issue of runtime and memory efficiency, each of which scales linearly with the number of models in the ensemble. Mixture-of-Experts (MoE) routines reduce the inference cost associated with ensembling to that of a single model (104; 89). Rather than averaging the predictions of each ensemble member, these methods learn to compose components—at the granularity of whole networks, layers, or individual neurons—from each model to compute the final output prediction. This amounts to learning how to construct a new neural network for each input presented. Traditional MoE learns a routing network which then selects $k$ of the ensemble members (“experts”) and computes a linear combination of their outputs to form the final prediction logits. This routing procedure was simplified by Fedus et al. (33) to produce Switch Transformers which demonstrate that even using just a single expert per input token is sufficient for strong performance.
3.3. Limitations of Ensembling
@@
Ensembles of several models have become much more straightforward to obtain thanks to advancements in our understanding of the loss landscape (41; 9; 52). Still, ensemble techniques suffer from poor scaling in terms of runtime and memory costs, as we must use multiple models for inference. MoE methods are a boon for use cases where inference cost is a limiting factor; however, they also suffer from several drawbacks. While the effective parameter count for a given input in an MoE model is lower, the total number of parameters in the mixture may be the same or even greater (98). As individual “experts” in the mixture specialize in just a few prediction domains, we must train more parameters in total to achieve strong performance across all domains. Consequently, training and serving MoE models requires constantly exchanging large numbers of parameters from memory. This requires a significant investment in complex infrastructure to serve models of billions or trillions of parameters. MoE models also exhibit decreased sample efficiency during training: each parameter is not trained during every batch, resulting in increased training costs. Though these methods offer exciting avenues for scaling and deploying ensemble models more efficiently, model merging offers the potential for equivalent performance at a significant cost reduction.
@@
Figure 3. Mode Convexity: Within an objective basin, one can linearly interpolate between known solutions to discover equally performant models with diverse behaviors.
4. Weight Aggregation
Table 2. Overview of Weight Aggregation techniques
Technique Simple Average Weighted Average Requires Data Requires Labels
FedSGD (91)
Stochastic Weighted Averaging (62)
Model Soup (124)
Sparse Soup (144)
SWAMP (20)
Task Arithmetic (58)
Model Breadcrumbs (24)
TIES-merging (130)
DARE (134)
FedAvg (109)
AdaMerging (133)
RegMean (64)
ZipIt! (108)
Evolutionary Model Merging (4)
Fisher Weighted Average (90)
MaTS (110)
@@ 4. Weight Aggregation
The fundamental technique underlying most model merging methods is weight aggregation. Given two or more models of the same architecture, one can compute an average or convex combination of the values for each corresponding parameter to obtain a set of new parameter values that comprise the merged model. These techniques are predicated on the principles of Bayesian Model Averaging, seeking to reduce estimation error from individual model bias and variance by combining several models for inference (48). Though weight aggregation methods show promise compared to ensemble methods, they also introduce several challenges. Weight aggregation methods usually introduce additional hyperparameters that must be set via trial-and-error or by some heuristic. Additionally, their limitations are not fully understood: while some characterizations of their failure cases can be extrapolated from the literature (see Section 5 below), more context on loss landscape geometry is necessary to fully understand their modes of failure. Below, we discuss several modes of weight aggregation, including simple averaging, weight steering via task arithmetic, data-aware weightings, and information-theoretic weightings. Table 2 lists the weight aggregation techniques we cover in this section and enumerates some of their distinguishing features.
4.1. Simple Averaging
@@
The most straightforward method of merging model parameters is to take a simple average over each corresponding parameter between two or more models, see Eq. (missing) 3. This technique is also employed in federated settings as FedSGD (91). However, simple averaging does not consider each model’s relevance to the task at hand. This method also consistently fails if the training distributions or the training trajectories between the two models have diverged (107; 64; 3). Despite this limitation, simple averaging is particularly useful in the “pre-train and fine-tune” model deployment paradigm, where models naturally share an architecture and are predisposed towards similar representations through pre-training (142; 141; 31). LLMs are a good example of this kind of deployment.
(3) $\theta^{\star}=\frac{1}{N}\sum_{i}\theta_{i}$
@@
A host of methods attempt to improve performance by using variations of simple averaging. Stochastic Weight Averaging (SWA) (62) averages checkpoints from a single model at different points in its training trajectory, demonstrating comparable performance to FGE (41). Model Soups (124) are simple averages of fine-tuned models of the same architecture trained with different hyperparameters. Model soups often enhance Out-Of-Distribution (OOD) and zero-shot performance on downstream tasks. Sparse soups (144) reduce the inference costs associated with model soup methods by enforcing a sparsity constraint via an iterative prune-retrain fine-tuning technique inspired by Iterative Magnitude Pruning (IMP) (113). Similarly, the SWAMP (20) method iteratively performs IMP and simple weight averaging between multiple fine-tuned models to produce a final sparse model. These works support contemporary observations that models closer to the center of a group of fine-tuned models stemming from the same pre-trained model are more generalizable and robust (44; 63). Wortsman et al. (123) demonstrate that several networks can be trained in tandem such that they form a linearly connected subspace of solutions. The midpoint of this subspace approaches the accuracy of an ensemble of independently trained models and demonstrates robustness to label noise. Model Stocks (63) take advantage of this observation by estimating a “central” network using only two fine-tuned models, significantly reducing the training effort required to construct model soups.
4.2. Weight Steering
@@
Task Arithmetic (58) is a technique for manipulating model weights to enhance task-specific and OOD performance. A task vector—defined as the difference between the pre-trained and fine-tuned model weights for a task, see Eq. (missing) 4—can be thought of as the direction in weight space that the pre-trained model must travel to achieve better accuracy on that particular task.
(4) $\Delta_{\text{task}}=\theta_{\text{task}}-\theta_{\text{pre}}$
@@
Interestingly, combining task vectors via simple arithmetic operations (addition and subtraction) provides predictable control over model behavior. For example, subtracting a task vector from the pre-trained weights reduces performance, while adding the same task vector improves performance on the corresponding task; adding two task vectors for distinct tasks to the pre-trained model weights enhances the resulting model’s performance on both tasks. This property is exploited by Model Breadcrumbs (24), a method that takes task vectors from several fine-tuned models, masks outlier directions, and combines them via addition into a single parameter update for the pre-trained checkpoint. Steering weights using task vectors can enable good performance in multiple tasks without degrading performance in any individual task significantly, particularly in larger models (59). TIES-merging (130) and DARE (134) improve upon Task Arithmetic by eliminating redundancies between task vectors, resulting in improved performance—particularly in the multi-task regime. It is worth noting the connection between these methods and parameter-efficient fine-tuning schemes such as LoRA (51; 45; 85) and IA3 (84). LoRA trains low-rank decompositions of weight matrices during fine-tuning that are added to the model weights for inference, while IA3 trains per-task scaling vectors in each layer that rescale the activations in that layer—equivalent to rescaling the weight matrix along corresponding directions in weight space. It is common to “swap” fine-tuning weights for different tasks by adding and subtracting out LoRA weights—a procedure analogous to how we use task vectors. LoReFT (125), inspired by LoRA, uses Distributed Alignment Search (DAS) (42; 126) to find task-specific directions in which to update low-rank fine-tuning weights.
@@
The success of weight steering methods implies that, within a single objective basin, we can meaningfully assign semantic properties to directions in weight space. Furthermore, we can leverage the linearly connected nature of local minima to smoothly interpolate between task-optimal parameter settings within a basin—even for multiple tasks (114). This raises the question: can we improve the performance of model merging techniques by calibrating the contribution of task-specific parameters to the final merged model?
@@
Figure 4. RegMean finds a new set of merged parameters that closely approximates the activation maps of the given source models in a manner analogous to least-squares regression.
4.3. Data-Aware Weightings